# Thermodynamic Descriptors from Molecular Dynamics as Machine Learning Features for Extrapolable Property Prediction

### 12 Mar 2026 v1

@@ III.3 MD-based models demonstrate superior extrapolation to structurally novel and complex chemical systems
Given the superior performance of the models trained on OPLS4-derived features, as demonstrated in the previous section, we selected these specific models for the subsequent extrapolation and benchmarking tests. A critical measure of a machine learning model’s utility is not just its accuracy within its training domain, but its ability to generalize to new, structurally distinct chemical matter. This capability is of crucial relevance for industrial applications in the pharmaceutical, crop science, and chemical sectors, where the exploration of novel molecular structures is essential for generating intellectual property and discovering materials with unprecedented properties. To evaluate the extrapolative power of our approach, we benchmarked our models against several well-established boiling point predictors: the group-contribution-based Joback [49] and Rarey-Nannoolal [75] methods, as well as a state-of-the-art graph neural network, GRAPPA [44]. A key factor in this selection was their practical accessibility via public web tools that provide instantaneous predictions directly from a SMILES string, such as those for Joback [15] and GRAPPA [45], allowing for a reproducible benchmark. The benchmark was performed on our rigorous external benchmark set (see section S.II.A). This set was deliberately curated to probe the limits of extrapolation and consists of 32 structurally diverse active pharmaceutical ingredients with varied complexity sourced from the DrugBank database [109]. These molecules were chosen to represent the kind of structurally complex, real-world chemical matter encountered in industrial discovery pipelines. Crucially, while complex, they are composed of standard organic elements, placing them squarely within the theoretical applicability domain of the chemoinformatic predictors we benchmark. This design ensures a direct and fair head-to-head comparison of true extrapolative performance, with the data also being curated to ensure high-quality experimental values and compatibility across all models. It is important to note the practical trade-off between predictive speed and extrapolative power. While chemoinformatic models are instantaneous, our physics-augmented approach requires an upfront simulation step. However, this computational cost is modest and highly accessible. On a modern GPU, a 20 ns simulation for a single compound is typically completed in 3-4 hours. Even on standard CPUs, the task takes approximately 15-20 hours using only 2 cores. This makes it feasible to generate descriptors for dozens of compounds simultaneously on a modern multi-core workstation or a small high-performance computing cluster, representing a practical and manageable investment to unlock the model’s robust extrapolative capabilities.
@@
Figure 4A plots the Mean Absolute Error (MAE) in Kelvin for all models on this challenging extrapolation test set. The corresponding Root Mean Squared Error (RMSE) plot, which reinforces the same trends, is provided in the Supplementary Information (Fig.S2.A). A surface-level comparison of the MAE suggests that the literature models GRAPPA and Rarey-Nannoolal achieve the highest accuracy. However, this comparison requires careful contextualization. GRAPPA’s high performance is likely attributable to its strength in interpolation, given its vast training set of  20,000 compounds. A Tanimoto similarity analysis supports this view. While the Tanimoto coefficient is a simple heuristic for molecular similarity, in this case it effectively delineates the domain of applicability. The test set compounds exhibit a high average similarity of 0.82 to GRAPPA’s training set, but a low similarity (0.38) to our own. This confirms that for our models, this benchmark represents a significantly more challenging ”out-of-distribution” extrapolation task.
@@
Importantly, this out-of-distribution challenge reveals a complete reversal of the performance hierarchy observed during interpolation. As seen in Figure 4A, our MD-only model now clearly outperforms the other models trained in this work, with an MAE of 31.0K. This result highlights a known limitation of chemoinformatics descriptors: features that are effective for interpolating within a known chemical space become unreliable when the model is faced with novel structures containing fragments or motifs absent from the training data [102]. The performance degradation is systematic: the chemoinformatics-only model fails (MAE = 53.5K), and the hybrid model (MAE = 40.4K) is penalized by the inclusion of these same unreliable structural features. In its own right, a model like GRAPPA provides fast, high-precision predictions and remains the superior tool within its well-defined domain. However, our results clearly show that for true out-of-distribution extrapolation, a model anchored in the underlying physics is fundamentally more robust. This finding aligns with recent conclusions in other domains, such as protein-ligand binding affinity prediction [59].
@@
The success of GRAPPA [44] in the initial benchmark demanded a more granular, head-to-head comparison to dissect the true generalization capabilities of both our MD-only model and the GNN. To achieve this, we stratified the external test set by Tanimoto similarity, measuring how structurally similar each test compound is to its nearest neighbor in each model’s respective training set (a full breakdown is provided in S.I of Supplementary Information). This analysis (Fig.S2.B) reveals their fundamentally different strengths. In the high-similarity regime (Tanimoto ¿ 0.95), GRAPPA’s performance is exceptional, with its MAE dropping to 4.1K for compounds most similar to its training data, confirming its power as a high-precision interpolator. To assess true extrapolative capability, we therefore focused the main analysis on the more challenging similarity range below 0.9 (Fig. 4B). In this view, a clear and consistent trend emerges where our MD-only model maintains a superior performance. For compounds with moderate similarity (0.75–0.9), our model’s MAE is 6.4K versus 9.8K for GRAPPA. This performance gap widens substantially as compound novelty increases, with our model achieving an MAE of 12.6K versus 26.9K in the medium similarity range, and 28.3K versus 40.9K for the most dissimilar compounds.
@@
The contrast in robustness becomes even more stark when normalizing the performance against each model’s own optimal baseline. GRAPPA’s error escalates from its 4.1K interpolation baseline by a factor of approximately 10 to predict the most dissimilar compounds. In stark contrast, our MD-only model exhibits a far more controlled response, with its MAE growing from its own 6.4K baseline by a factor of only 4.4 in the same low-similarity regime. This analysis confirms their distinct predictive behaviors: while GRAPPA excels at high-precision interpolation, our physics-informed model demonstrates a more graceful and predictable degradation of performance when faced with structural novelty. These results validate the physics-augmented strategy as a robust foundation for building more reliable predictive tools, particularly for the challenging low-similarity regime where traditional models falter.
@@
The ultimate advantage of a physics-based approach is the ability to predict properties for chemical classes that lie entirely outside the domain of traditional models. Many existing predictors, including GRAPPA, are restricted to neutral organic molecules containing a limited set of elements (typically C, H, O, N, S, P, and sometimes various halogens) and cannot handle charged species or exotic elements. Our MD-based framework has no such intrinsic limitations. Fig. 4C showcases the model’s capability to extrapolate to these systems even though they were not in the training set. The first group (pink circles) includes neutral compounds containing uncommon elements such as Si, B, and Te, or molecules lacking carbon altogether (e.g., tribromosilane). The second group (blue stars) comprises charged systems, including salts (e.g., Acesulfame K), ionic liquids (ILs), and ternary mixtures containing ILs and solvents. We note that compiling this validation set required extensive literature curation, as reliable, experimentally measured boiling points for such systems are scarce, particularly for salts and ILs which often decompose at high temperatures [3]. The ability of our model to make reasonable predictions for these complex systems, which are inaccessible to many widely used tools, underscores the power of leveraging fundamental thermodynamic properties for molecular property prediction and opens a path toward a more universal predictive framework.
@@ IV Conclusions
In this work, we have developed a physics-augmented machine learning framework that demonstrates how first-principles-informed networks can perform more robustly when extrapolating. While existing methods like the group contribution-based Rarey-Nannoolal [75] or machine learning models like GRAPPA [44] are competitive, our work highlights a specific strength in navigating structural novelty. Our approach successfully integrates the regression power of modern machine learning with the fundamental principles of thermodynamics and explicit MD simulations, offering a complementary tool for exploring the novel chemical spaces.
Our first critical step was to validate the physical premise of our approach, establishing a strong, direct correlation between simulated cohesive energies and experimental boiling points across multiple force fields and temperatures (Fig. 2). Building on this, we demonstrated that a minimal set of just three simulation-derived features—principally the heat of vaporization—is sufficient to build an ”MD-only” model with predictive accuracy competitive with complex chemoinformatics models that rely on thousands of abstract descriptors (Fig. 3B). This constitutes a feature dimensionality reduction of over two orders of magnitude, transforming the prediction from a ”black-box” exercise into an interpretable, physically-grounded regression. The model’s reliance on $\Delta H_{\text{vap}}$ confirms it learns the correct underlying physics, a level of transparency that purely structural models lack.
The primary strength of our framework is revealed when venturing beyond familiar chemical territory. Our physics-informed model’s performance is substantially less sensitive to growing structural dissimilarity, a stark contrast to the brittleness of purely data-driven approaches (Fig. 4A-B). This robustness establishes our model as a reliable complementary tool, capable of predicting boiling points for systems inaccessible to most established methods, including inorganic compounds, molecules with uncommon elements like silicon and boron, and charged systems such as ionic liquids (Fig. 4C). By anchoring predictions to fundamental thermodynamics, we demonstrate a clear path toward more generalizable ML models capable of navigating uncharted chemical space.
While this study is focused solely on normal boiling point predictions, our results provide strong, systematic evidence for a broader principle: MD-informed machine learning models are more robust for predicting condensed-phase properties governed by intermolecular forces. The framework presented here is not a single-purpose solution but a promising research direction. Rooting predictions in thermodynamics offers a pathway to develop more accurate and widely applicable models, essential for navigating the unexplored chemical territories vital to the future of materials science, pharmacology, drug discovery and engineering.
@@ V Acknowledgements
N.H.E. acknowledges funding from Bayer Aktiengesellschaft, Division CropScience residing at Kaiser-Wilhelm-Allee 1, 51373 Leverkusen, Germany (“Bayer”). J.R. acknowledges funding from the Spanish Ministry of Economy and Competitivity (PID2022-136919NB-C32). J.R.E. acknowledges funding from Emmanuel College, the University of Cambridge, the Ramon y Cajal fellowship (RYC2021-030937-I), the Spanish scientific plan and committee for research reference PID2022-136919NA-C33, and the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (grant agree- ment no. 101160499). This work has been performed using resources provided by the HPC system in Bayer. The authors also thankfully acknowledge RES computational resources provided by Mare Nostrum 5 through the activity 2024-3-0001. Finally, the authors thank Marco Hoffmann for generously providing GRAPPA’s training set data, which enabled a direct and valuable comparison between models.

### 18 Mar 2026 v2

@@ III.3 MD-based models demonstrate superior extrapolation to structurally novel and complex chemical systems
Given the superior performance of the models trained on OPLS4-derived features, as demonstrated in the previous section, we selected these specific models for the subsequent extrapolation and benchmarking tests. A critical measure of a machine learning model’s utility is not just its accuracy within its training domain, but its ability to generalize to new, structurally distinct chemical matter. This capability is of crucial relevance for industrial applications in the pharmaceutical, crop science, and chemical sectors, where the exploration of novel molecular structures is essential for generating intellectual property and discovering materials with unprecedented properties. To evaluate the extrapolative power of our approach, we benchmarked our models against several well-established boiling point predictors: the group-contribution-based Joback [50] and Rarey-Nannoolal [76] methods, as well as a state-of-the-art graph neural network, GRAPPA [45]. A key factor in this selection was their practical accessibility via public web tools that provide instantaneous predictions directly from a SMILES string, such as those for Joback [15] and GRAPPA [46], allowing for a reproducible benchmark. The benchmark was performed on our rigorous external benchmark set (see section S.II.A). This set was deliberately curated to probe the limits of extrapolation and consists of 32 structurally diverse active pharmaceutical ingredients with varied complexity sourced from the DrugBank database [111]. These molecules were chosen to represent the kind of structurally complex, real-world chemical matter encountered in industrial discovery pipelines. Crucially, while complex, they are composed of standard organic elements, placing them squarely within the theoretical applicability domain of the chemoinformatic predictors we benchmark. This design ensures a direct and fair head-to-head comparison of true extrapolative performance, with the data also being curated to ensure high-quality experimental values and compatibility across all models. It is important to note the practical trade-off between predictive speed and extrapolative power. While chemoinformatic models are instantaneous, our physics-augmented approach requires an upfront simulation step. However, this computational cost is modest and highly accessible. On a modern GPU, a 20 ns simulation for a single compound is typically completed in half an hour. This makes it feasible to generate descriptors for dozens of compounds simultaneously on a modern multi-core workstation or a small high-performance computing cluster, representing a practical and manageable investment to unlock the model’s robust extrapolative capabilities.
@@
Figure 4A plots the Mean Absolute Error (MAE) in Kelvin for all models on this challenging extrapolation test set. The corresponding Root Mean Squared Error (RMSE) plot, which reinforces the same trends, is provided in the Supplementary Information (Fig.S2.A). A surface-level comparison of the MAE suggests that the literature models GRAPPA and Rarey-Nannoolal achieve the highest accuracy on this dataset.
@@
However, this simple comparison is misleading. A deeper analysis reveals a lower MAE for our thermodynamics-based models once the nature of the extrapolation task is considered. The test set compounds exhibit a high average Tanimoto similarity of 0.82 to GRAPPA’s training set, but a low similarity (0.38) to our own. This confirms that for our models, trained on only  1,000 compounds versus GRAPPA’s 20,000, this benchmark represents a significantly more challenging out-of-distribution extrapolation task. In this true extrapolative regime, our MD-only model clearly outperforms the other models trained in this work, achieving an MAE of 31.0 K 4A. This result highlights a known limitation of chemoinformatics descriptors: features that are effective for interpolating within a known chemical space become unreliable when the model is faced with novel structures containing fragments or motifs absent from the training data [104]. The performance degradation is systematic: our chemoinformatics-only model fails (MAE = 53.5 K), and the hybrid model (MAE = 40.4 K) is penalized by including these same unreliable structural features. While a state-of-the-art GNN like GRAPPA does not use traditional chemoinformatics descriptors, its performance is also tightly coupled to training data similarity. Therefore, while GRAPPA remains the superior tool for high-precision predictions within its vast, well-defined domain, our results demonstrate that for true out-of-distribution extrapolation—especially when training data is limited—a model anchored in underlying physics is fundamentally more robust. This finding aligns with recent conclusions in other domains, such as protein-ligand binding affinity prediction [60].
@@
To dissect the true generalization capabilities of both our MD-only model and the GNN we stratified the external test set by Tanimoto similarity, measuring how structurally similar each test compound is to its nearest neighbor in each model’s respective training set (a full breakdown is provided in S.I of Supplementary Information). This analysis (Fig.S2.B) reveals their fundamentally different strengths. In the high-similarity regime (Tanimoto >> 0.75), GRAPPA’s performance is exceptional, with its MAE dropping to 4.1K for compounds most similar to its training data, confirming its power as a high-precision interpolator. To assess true extrapolative capability, we therefore focused the main analysis on the more challenging similarity range below 0.9 (Fig. 4B). In this view, a clear and consistent trend emerges where our MD-only model maintains a superior performance. For compounds with moderate similarity (0.75–0.9), our model’s MAE is 6.4K versus 9.8K for GRAPPA. This performance gap widens substantially as compound novelty increases, with our model achieving an MAE of 12.6K versus 26.9K in the medium similarity range, and 28.3K versus 40.9K for the most dissimilar compounds.
@@
The disparity in robustness is most evident when considering normalized error growth. When extrapolating to the most dissimilar compounds, GRAPPA’s error escalates tenfold from its 4.1 K interpolation baseline. In contrast, our MD-only model exhibits a far more controlled response, its error growing by a factor of only 4.4 from its 6.4 K baseline. This confirms their distinct behaviours. While GRAPPA excels at high-precision interpolation, our physics-augmented approach provides a more graceful degradation in performance, establishing it as a more robust foundation for prediction in low-similarity regimes.
@@
The ultimate advantage of a physics-based approach is the ability to predict properties for chemical classes that lie entirely outside the domain of traditional models. Existing predictors, including GRAPPA, are restricted to neutral organic molecules containing a limited set of elements (typically C, H, O, N, S, P, and sometimes various halogens) and cannot handle charged species or exotic elements. Fig. 4C showcases the model’s capability to extrapolate to these systems even though they were not in the training set. The first group (pink circles) includes neutral compounds containing uncommon elements such as Si, B, and Te, or molecules lacking carbon altogether (e.g., tribromosilane). The second group (blue stars) comprises charged systems, including salts (e.g., Acesulfame K), ionic liquids (ILs), and ternary mixtures containing ILs and solvents. We note that experimentally measured boiling points for such systems are scarce, particularly for salts and ILs which often decompose at high temperatures [3]. The ability of our model to make reasonable predictions for these complex systems, which are inaccessible to many widely used tools, underscores the power of leveraging fundamental thermodynamic properties for molecular property prediction and opens a path toward a more universal predictive framework.
@@ IV Conclusions
In this work, we have developed a physics-augmented machine learning framework that demonstrates how first-principles-informed networks can perform more robustly when extrapolating. While existing methods like the group contribution-based Rarey-Nannoolal [76] or machine learning models like GRAPPA [45] are very competitive around the training set, our work highlights a specific strength in navigating structural novelty. Our approach successfully integrates the regression power of modern machine learning with the fundamental principles of thermodynamics and explicit MD simulations, offering a reliable tool for exploring the novel chemical spaces.
We demonstrated that a minimal set of three simulation-derived features yields a model with predictive accuracy competitive with complex chemoinformatics models that rely on thousands of abstract descriptors (Fig. 3B). The resulting framework is, moreover, interpretable and physically grounded. This approach successfully integrates the regression power of modern machine learning with the fundamental principles of thermodynamics, offering a reliable tool for exploring novel chemical spaces.
The primary strength of our framework is revealed when venturing beyond chemical space of the training set. Our physics-informed model’s performance is substantially less sensitive to structural dissimilarity, in contrast to purely data-driven approaches (Fig. 4A-B). This robustness establishes our model as a reliable tool, capable of predicting boiling points for systems inaccessible to most established methods, including inorganic compounds, molecules with uncommon elements like silicon and boron, and charged systems such as ionic liquids (Fig. 4C). By anchoring predictions to molecular thermodynamics, we demonstrate a clear path toward more generalizable ML models capable of navigating uncharted chemical space.
While this study is focused solely on normal boiling point predictions, our results provide strong, systematic evidence for a broader principle: MD-informed machine learning models are more robust for predicting properties governed by intermolecular forces. Rooting predictions in thermodynamics offers a pathway to develop more accurate and widely applicable models, essential for navigating the unexplored chemical territories vital to the future of materials science, pharmacology, drug discovery and engineering.
@@ V Acknowledgements
N.H.E. acknowledges funding from Bayer Aktiengesellschaft at Kaiser-Wilhelm-Allee 1, 51373 Leverkusen, Germany (“Bayer”). J.R.E. acknowledges funding from Emmanuel College, the University of Cambridge, the Ramon y Cajal fellowship (RYC2021-030937-I), the Spanish scientific plan and committee for research reference PID2022-136919NA-C33, and the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (grant agree- ment no. 101160499). J.R. acknowledges funding from the Spanish Ministry of Economy and Competitivity (PID2022-136919NB-C32). This work has been performed using resources provided by the HPC system in Bayer. The authors also thankfully acknowledge RES computational resources provided by Mare Nostrum 5 through the activity 2024-3-0001. Finally, the authors thank Marco Hoffmann for generously providing GRAPPA’s training set data, which enabled a direct and valuable comparison between models.